# Targeted Immunisation Thresholds for the Contact Process on Power-Law Trees

### v1

@@ 4.2 Subcritical truncation threshold
will have, given $\lambda>0$ sufficiently small, $P_{\rm{surv}}^{\mathbbm{t}}(\lambda)=0$. For $\tau=2.5$ and some $\rho$ with
$\rho\leq\frac{1}{4},\quad\frac{2c_{q}\sqrt{\rho}}{1+\sqrt{1-4\rho}}<1$
@@
The infection process dies out a.s. if we can show that the probability of the infection reaching distance $k$ from the root tends to 0 as $k\to\infty$. To this end, we simply combine the mean number of vertices at depth $k$ given by (7) with the probability of infecting any given such vertex upper bounded by Lemma 2.1. For $\tau=2.5$, we take $\delta=\rho$ in Lemma 2.1 to obtain
$\frac{2\lambda\operatorname{\mathbb{E}}[D]}{1+\sqrt{1-4\rho}}\to\frac{2c_{q}\sqrt{\rho}}{1+\sqrt{1-4\rho}}<1,$
which taken to the power $k$ indeed tends to 0. For $\tau<2.5$, the truncation threshold is $o(\lambda^{-2})$, and thus bounded by $\delta/\lambda^{2}$ with arbitrary $\delta>0$. We conclude in that case by observing
$\frac{2\lambda\operatorname{\mathbb{E}}[D]}{1+\sqrt{1-4\delta}}\to\frac{2}{1+\sqrt{1-4\delta}}c_{q}\frac{\tau-2}{3-\tau}\rho^{3-\tau},$
@@ 5.1 Large truncation threshold
When $2.5<\tau\leq 3$ we will ensure survival of the infection by using again a SIR-like lower bound, with the refinement that big vertices use their small children to maintain the infection locally for a long enough time so as to typically infect its big children.
@@
More precisely, say a vertex is big if it has degree larger than $\mathbbm{s}=\mbox{$\frac{1}{\varepsilon}$}\lambda^{-2}\log(1/\lambda)$, where $\varepsilon$ is as in Lemma A.1, and assume $\rho>1/\varepsilon$, so that the truncation threshold $\mathbbm{t}=\rho\lambda^{-2}\log(1/\lambda)$ satisfies $\mathbbm{t}\geq\mathbbm{s}$. If a vertex $x$ is big and has $D_{x}\geq\mathbbm{s}$ children, we can use Lemma A.1 to argue that with probability at least $1-\varepsilon c/\log(1/\lambda)$, which is larger than $1/2$ for small $\lambda$, the infection survives at $x$ for a total time at least $\exp(\log(1/\lambda))=1/\lambda$, using only infections going back and forth to its first $\mathbbm{s}$ children. Write $L_{x}$ for this event, for “local survival” at $x$. Conditionally on $L_{x}$, the remaining $D_{x}-\mathbbm{s}$ children are independently infected with probability at least $21-e^{-1}\geq 1/$ 2. Thus the number of (remaining) infected vertices is bounded below by the mixed binomial offspring distribution
$Y\big{|}D_{x},\mathbbm{1}_{L_{x}}\sim\operatorname{Bin}\left((D_{x}-\mathbbm{s})\mathbbm{1}_{D_{x}\geq\mathbbm{s}}\mathbbm{1}_{L_{x}},\frac{1}{2}\right).$
We then consider each of these infected children as a source of an independent infection living in its own branch for which we use the same survival strategy, revealing a Galton-Watson process with offspring distribution $Y$, whose survival implies survival of the contact process. As for the case $\tau\leq 2.5$, we lower bound the survival probability by a second moment method on $Y$ using Lemma 4.3. We have
$\operatorname{\mathbb{E}}[Y]\geq\frac{1}{4}\operatorname{\mathbb{E}}[(D_{x}-\mathbbm{s})\mathbbm{1}_{D_{x}\geq\mathbbm{s}}]\begin{cases}=\Omega(\mathbbm{t}^{3-\tau})&\mbox{ for }\tau<3\\ \to\frac{c_{q}}{6}\left(\log(\varepsilon\rho)-1+\frac{2}{\varepsilon\rho}\right)&\mbox{ for }\tau=3\end{cases}$
In the case $\tau=3$ we further suppose $\rho$ large enough so as to ensure $\operatorname{\mathbb{E}}[Y]>1$. The variance of $Y$ has
$\operatorname{Var}(Y)$ $=\operatorname{\mathbb{E}}\left[\operatorname{Var}(Y\big{|}D_{x},1_{L_{x}})\right]+\operatorname{Var}\left(\operatorname{\mathbb{E}}[Y\big{|}D_{x},1_{L_{x}}]\right)$
$\leq\operatorname{\mathbb{E}}\left[\frac{1}{4}(D_{x}-\mathbbm{s})\mathbbm{1}_{D_{x}\geq\mathbbm{s}}\mathbbm{1}_{L_{x}}\right]+\operatorname{Var}\left(\frac{1}{2}(D-S)\mathbbm{1}_{D\geq S}\mathbbm{1}_{A}\right)$
@@
where in the second term we bound the variance by the expectation of the square of the random variable and use that that random variable is bounded by $(\mathbbm{t}-\mathbbm{s})/2$. By Lemma 4.3 we then obtain a lower bound for the survival probability of order
@@ 5.2 Small truncation threshold
which, although being large, will be smaller than $\lambda^{-1}$ by the choice of $\rho$. Thus even a vertex of maximal degree has only small probability of ever infecting any given direct neighbour, which one might see as a first indication that getting upper bounds in that case might be feasible.
To provide a rigorous upper-bound, though, we will use the Contact process-Branching random walk of Section 2.3, with sinks all “big vertices”, namely all vertices with degree larger than $1/8\lambda^{2}$.
We first consider a single big vertex and let $(\mathscr{T},o)$ be the infinite rooted tree in which
$\operatorname{d}(o)=\mathbbm{t},\qquad\forall v\in V(\mathscr{T})\setminus\{o\}\quad\operatorname{d}(v)=\left\lfloor\frac{1}{8\lambda^{2}}\right\rfloor.$
@@
On the tree $(\mathscr{T},o)$ we have
$\mathbb{E}\left(\int_{0}^{\infty}\xi_{t}(o){\rm d}t\Bigg{|}\xi_{0}=\mathbbm{1}_{o}\right)\leq 88\lambda^{-72\rho}$
@@
We wait for this sum to go below level $8\lambda\mathbbm{t}$ and for the root of the tree to recover, and then check if this is a permanent recovery for the root vertex. If not, we repeat the procedure up until a successful recovery, providing in each step bounds on the expected time taken by the step, as well as on the probability of the recovery to be successful.
@@
From the calculation in [mountford2013metastable] we have the bound
$\frac{\mathbb{E}\left({\rm d}H_{t}\right)}{{\rm d}t}\leq\lambda\mathbbm{t}\,\xi_{t}(o)-\frac{1}{4}H_{t}$ (8)
noting also that if $\xi_{t}(o)=0$, because neighbours of the root have weight $1$, then this root is reinfected at rate bounded by $\lambda H_{t}$. Define further
$Y_{t}:=e^{t/4}\left(H_{t}-4\lambda\mathbbm{t}\right)$
which we find is a supermartingale by the product rule, using of course that $\xi_{t}(o)\leq 1$.
$H$ has no upper limit, but in the event that it is large we can bound the expected hitting times for $H$ then by Markov’s inequality on $Y_{t}+4\lambda\mathbbm{t}e^{t/4}$:
$\begin{split}\mathbb{P}\left(H_{t}\geq 8\lambda\mathbbm{t}\Big{|}H_{0}\right)&=\mathbb{P}\left(Y_{t}+4\lambda\mathbbm{t}e^{t/4}\geq 8\lambda\mathbbm{t}e^{t/4}\Big{|}Y_{0}\right)\\ &\leq\frac{\mathbb{E}\left(Y_{t}+4\lambda\mathbbm{t}e^{t/4}\Big{|}Y_{0}\right)}{8\lambda\mathbbm{t}e^{t/4}}\leq\frac{Y_{0}}{8\lambda\mathbbm{t}e^{t/4}}+\frac{1}{2}.\end{split}$
On the event $H_{0}>8\lambda\mathbbm{t}$, we can pick $t_{1}=t_{1}(Y_{0})\leq 4\log\left(\frac{Y_{0}}{2\lambda\mathbbm{t}}\right)$ such that we have exactly $\frac{1}{4}$ chance of upper bounding $H_{t_{1}}$. If we fail:
$\mathbb{E}\left(Y_{t_{1}}\Big{|}Y_{0},\left\{H_{t_{1}}\geq 8\lambda\mathbbm{t}\right\}\right)\leq\frac{\mathbb{E}\left(Y_{t_{1}}\Big{|}Y_{0}\right)}{\mathbb{P}\left(H_{t_{1}}\geq 8\lambda\mathbbm{t}\right)}=\frac{\mathbb{E}\left(Y_{t_{1}}\Big{|}Y_{0}\right)}{3/4}\leq\frac{4}{3}Y_{0}$
and thus we can bound the mean of $Y$ over multiple attempts to go below $8\lambda\mathbbm{t}$. Suppose we have constructed times $t_{1},\dots,t_{k-1}$ with
$\forall i\leq k-1,\quad\mathbb{P}\left(H_{t_{i}}\leq 8\lambda\mathbbm{t}\Big{|}Y_{0},\left\{\min_{j<i}H_{t_{j}}\geq 8\lambda\mathbbm{t}\right\}\right)=\frac{1}{4}$
so to get another $\frac{1}{4}$ chance of bounding $H_{t_{k}}\leq 8\lambda\mathbbm{t}$ we have to wait for a time $t_{k}$ with
$\begin{split}\mathbb{E}\left(t_{k}-t_{k-1}\Big{|}Y_{0},\left\{\min_{j<k}H_{t_{j}}\geq 8\lambda\mathbbm{t}\right\}\right)&\leq\mathbb{E}\left(4\log\left(\frac{Y_{t_{k-1}}}{2\lambda\mathbbm{t}}\right)\Bigg{|}Y_{0},\left\{\min_{j<k}H_{t_{j}}\geq 8\lambda\mathbbm{t}\right\}\right)\\ &\leq 4\log\left(\mathbb{E}\left(\frac{Y_{t_{k-1}}}{2\lambda\mathbbm{t}}\Bigg{|}Y_{0},\left\{\min_{j<k}H_{t_{j}}\geq 8\lambda\mathbbm{t}\right\}\right)\right)\\ &\leq 4(k-1)\log\frac{4}{3}+4\log\left(\frac{Y_{0}}{2\lambda\mathbbm{t}}\right)\end{split}$
by Jensen’s inequality. So, iteratively, on the event $H_{0}>8\lambda\mathbbm{t}$, we can bound the expected hitting time of $[0,8\lambda\mathbbm{t}]$ given $H_{0}$ by
$\sum_{k=1}^{\infty}\left(\frac{3}{4}\right)^{k-1}\left(4(k-1)\log\frac{4}{3}+4\log\left(\frac{Y_{0}}{2\lambda\mathbbm{t}}\right)\right)=48\log\frac{4}{3}+16\log\left(\frac{H_{0}-4\lambda\mathbbm{t}}{2\lambda\mathbbm{t}}\right),$ (9)
Note we can further bound this by
$48\log\frac{4}{3}+16\log\left(1+\frac{H_{0}}{2\lambda\mathbbm{t}}\right),$
inequality that clearly stays satisfied on the event $H_{0}<8\lambda\mathbbm{t}$. So Jensen’s inequality again gives an expected hitting time of $[0,8\lambda\mathbbm{t}]$ bounded by
@@
$48\log\frac{4}{3}+16\log\left(1+\frac{\operatorname{\mathbb{E}}[H_{0}]}{2\lambda\mathbbm{t}}\right).$ (10)
Once we have hit this level at some time $t_{k}$, let $s\geq t_{k}$ be the next recovery time of $o$, or $s=t_{k}$ if $o$ is healthy at $t_{k}$. By (8) and the construction of $s$ we have $\mathbb{E}(H_{s}|H_{0})\leq 9\lambda\mathbbm{t}$.
Suppose that the centre is never reinfected. If this happens, by (8) we see that $(e^{t/4}M_{t})_{t>s}$ is a supermartingale. Hence by Markov’s inequality
$\mathbb{P}\left(\int_{s}^{\infty}H_{t}{\rm d}t>72\lambda\mathbbm{t}\Big{|}H_{0}\right)\leq\frac{9\lambda\mathbbm{t}\int_{s}^{\infty}e^{-(t-s)/4}{\rm d}t}{72\lambda\mathbbm{t}}=\frac{1}{2}$
and so, recalling that we infect the root at rate bounded by $\lambda M_{t}$, the probability that we really don’t ever see reinfection of the root is at least $\frac{1}{2}e^{-72\lambda^{2}\mathbbm{t}}\geq\frac{1}{2}\lambda^{72\rho}.$
If we fail this stage, and the centre is reinfected at some time $r>s$, then, using the supermartingale property of $(H_{(s+u)\wedge r})_{u\geq 0}$, we get
@@
$\mathbb{E}\left(H_{r}\Big{|}H_{0}\right)\leq\frac{9\lambda\mathbbm{t}}{1-\frac{1}{2}\lambda^{-72\rho}}\leq 10\lambda\mathbbm{t}$
which, by (10), gives an expected hitting time of $[0,8\lambda\mathbbm{t}]$ bounded by
$4348\log\frac{4}{3}+16\log 6<$ 43
and so the period until the next recovery attempt has expected length bounded by $44$, after considering the additional recovery period after hitting.
The initial condition we consider here actually has $H_{0}=0$ so we could get a minor improvement by separately considering the initial attempts, but we have at least shown by a geometric number of attempts that the total expected infectious period of the root is bounded by $\frac{44}{\frac{1}{2}\lambda^{72\rho}}$ as claimed. ∎
@@
Consider an infection sink on $\mathscr{T}$ at distance $\operatorname{d}(v,o)=k\geq 1$. For $\rho$ sufficiently small, the number of infections we expect it to receive from an infection at $o$ is
$m^{[v]}_{\infty}\leq 45\lambda^{-72\rho}(2\lambda)^{k}.$

### v2

@@ 4.2 Subcritical truncation threshold
For $\tau=2.5$ and some $\rho$ with
$\rho\leq\frac{1}{4},\quad\frac{2c_{X}\sqrt{\rho}}{1+\sqrt{1-4\rho}}<1$
@@
The infection process dies out a.s. if we can show that the probability of the infection reaching distance $k$ from the root tends to 0 as $k\to\infty$. To this end, we simply combine the mean number of vertices at depth $k$ given by (8) with the probability of infecting any given such vertex upper bounded by Lemma 2.2.
For $\tau=2.5$, we take $\delta=\rho$ in Lemma 2.2 to obtain
$\frac{2\lambda\operatorname{\mathbb{E}}[D]}{1+\sqrt{1-4\rho}}\to\frac{2c_{X}\sqrt{\rho}}{1+\sqrt{1-4\rho}}<1,$
which taken to the power $k$ indeed tends to 0.
For $\tau<2.5$, the truncation threshold is $o(\lambda^{-2})$, and thus bounded by $\delta/\lambda^{2}$ with arbitrary $\delta>0$. We conclude in that case by observing
$\frac{2\lambda\operatorname{\mathbb{E}}[D]}{1+\sqrt{1-4\delta}}\to\frac{2}{1+\sqrt{1-4\delta}}c_{X}\frac{\tau-2}{3-\tau}\rho^{3-\tau},$
@@
In this section we address the parameter regime $\tau\in(2.5,3]$ of our main theorem. First in Section 5.1 we lower bound the survival probability in the supercritical case, and finally in Theorem 5.5 we find our subcritical upper bound.
@@ 5.1 Large truncation threshold
When $2.5<\tau\leq 3$ we ensure survival of the infection by using again a SIR-like lower bound, with the refinement that big vertices use their small children to maintain the infection locally for a long enough time so as to typically infect its big children.
More precisely, say a vertex is big if it has offspring larger than $\mathbbm{s}=\mbox{$\frac{1}{\varepsilon}$}\lambda^{-2}\log(\nicefrac{{1}}{{\lambda}})$, where $\varepsilon$ is as in Lemma A.1. Note we will typically assume $\rho>\nicefrac{{1}}{{\varepsilon}}$, so that the truncation threshold $\mathbbm{t}=\rho\lambda^{-2}\log(\nicefrac{{1}}{{\lambda}})$ satisfies $\mathbbm{t}\geq\mathbbm{s}$, as is necessary to have any big vertex.
@@
Suppose the root vertex $o$ is big, with degree $D_{o}\geq\mathbbm{s}$, and consider the contact process evolving only on $o$ and its first $\mathbbm{s}$ children (namely we consider here only infections going back and forth to these $\mathbbm{s}$ children). If for this contact process the total time $o$ is infected is at least $\nicefrac{{1}}{{\lambda}}$, we say there is local survival at $o$, and write $L_{o}$ for this event. By Lemma A.1, the probability of $L_{o}$ is at least $1-\varepsilon c/\log(\nicefrac{{1}}{{\lambda}})$, which is larger than $\nicefrac{{1}}{{2}}$ for small $\lambda$.
Moreover, conditionally on $L_{o}$, the vertex $o$ is infected for a time period (at least) $\nicefrac{{1}}{{\lambda}}$, during which it communicates (at least) $\operatorname{Pois}(1)$ infections to each of its remaining $D_{o}-\mathbbm{s}$ children, which are therefore independently infected with probability at least $1-e^{-1}\geq\nicefrac{{1}}{{2}}$. We then consider each of these infected children as a source of an independent infection living in its own branch for which we use the same survival strategy.
Doing so, we can ensure survival of the contact process if we have survival of a Galton-Watson process with an offspring distribution $Y$ that is constructed in three steps:
  • First, pick the offspring of the vertex, $D\sim X\mathbbm{1}_{X\leq\mathbbm{t}}$.
  • If $D\geq\mathbbm{s}$, then pick $L\sim\operatorname{Ber}(\nicefrac{{1}}{{2}})$, where $L=1$ stems for local survival at the vertex.
  • Finally, pick $Y\big{|}D,L\sim\operatorname{Bin}\left(\mathbbm{1}_{D\geq\mathbbm{s}}L.(D-\mathbbm{s}),\frac{1}{2}\right),$ the number of the remaining $D-\mathbbm{s}$ children of the vertex to which the vertex communicates an infection.
As for the case $\tau\leq 2.5$, we lower bound the survival probability by a second moment method on $Y$ using Lemma 4.3. We have if $\tau<3$
$\operatorname{\mathbb{E}}[Y]=\frac{1}{2}\operatorname{\mathbb{E}}[\mathbbm{1}_{D\geq\mathbbm{s}}L.(D-\mathbbm{s})]=\frac{1}{4}\operatorname{\mathbb{E}}[(D-\mathbbm{s})\mathbbm{1}_{D\geq\mathbbm{s}}]=\Omega(\operatorname{\mathbb{E}}[D])=\Omega(\mathbbm{t}^{3-\tau})$
while in the case $\tau=3$:
$\operatorname{\mathbb{E}}[Y]=\frac{1}{4}\operatorname{\mathbb{E}}[(D-\mathbbm{s})\mathbbm{1}_{D\geq\mathbbm{s}}]=\frac{1}{4}\int_{\mathbbm{s}}^{\mathbbm{t}}\operatorname{\mathbb{P}}(D\geq z){\rm d}z\sim\frac{c_{X}}{4}\left(\log(\varepsilon\rho)-1+\frac{1}{\varepsilon\rho}\right)>1$
so that the final inequality is true whenever $\rho$ is large enough.
The variance of $Y$ has
$\operatorname{Var}(Y)$ $=\operatorname{\mathbb{E}}\left[\operatorname{Var}(Y\big{|}D,L)\right]+\operatorname{Var}\left(\operatorname{\mathbb{E}}[Y\big{|}D,L]\right)$
$\leq\operatorname{\mathbb{E}}\left[\frac{1}{4}\mathbbm{1}_{D\geq\mathbbm{s}}L.(D-\mathbbm{s})\right]+\operatorname{Var}\left(\frac{1}{2}\mathbbm{1}_{D\geq\mathbbm{s}}L.(D-\mathbbm{s})\right)$
@@
where in the second term we bound the variance by the expectation of the square of the random variable and use that the random variable is bounded by $(\mathbbm{t}-\mathbbm{s})/2$. By Lemma 4.3 we then obtain a lower bound for the survival probability of order
@@ 5.2 Small truncation threshold
which, although being large, will be smaller than $\lambda^{-1}$ by the choice of $\rho$. Thus even a vertex of maximal degree has only small probability of ever infecting any given direct neighbour, which one might see as a first indication that getting upper bounds in that case might be feasible. To provide a rigorous upper-bound, though, we will:
  • First consider the case of a single vertex of degree larger than $1/8\lambda^{2}$. In that case, we
use the Contact process-Branching random walk of Section 2.4, with sinks all “big vertices”, namely all vertices with degree larger than $1/8\lambda^{2}$.
As
@@
Consider $\left(\mathscr{T}_{\mathbbm{d},\left\lfloor\frac{1}{8\lambda^{2}}\right\rfloor},o\right)$ the infinite rooted tree with root $o$ of degree $\mathbbm{d}$ and all other vertices of degree $\left\lfloor\frac{1}{8\lambda^{2}}\right\rfloor$, and start the infection from only the root infected. Then the average total time the root is infected is bounded by
$\operatorname{\mathbb{E}}\left[\int_{0}^{\infty}\xi_{t}(o){\rm d}t\Big{|}\xi_{0}=\mathbbm{1}_{o}\right]\leq 53e^{72\lambda^{2}\mathbbm{d}}.$
@@
From the calculation [mountford2013metastable, Equation 5.1] we can bound its expected infinitesimal increase by
@@
$\frac{\operatorname{\mathbb{E}}\left[H_{t+\rm dt}-H_{t}|\mathcal{F}_{t}\right]}{{\rm d}t}\leq\lambda\mathbbm{d}\,\xi_{t}(o)-\frac{1}{4}H_{t},$ (9)
where $\mathcal{F}_{t}$ is the $\sigma-$algebra generated by $(H_{s})_{s\leq t}$. By these dynamics, the process $H_{t}$ is likely to quickly go to 0 when the root is healthy, while it is likely to quickly go to the value $4\lambda\mathbbm{d}$ (or close to this value) when the root is infected. Of course, the root also changes its state from infected to healthy at rate 1, and from healthy to infected at rate $\lambda$ times the number of infected neighbours, which is bounded above by $\lambda H_{t}$, as neighbours of the root have weight 1.
To bound the total time the root is infected, we first check whether its first recovery is a permanent recovery or not. If this fails, we first ask the process $H_{t}$ to go below level $8\lambda\mathbbm{d}$ before checking for the next recovery and repeating the procedure.
More precisely, we introduce the sequences of stopping times $(R_{k})_{k\geq 1}$, $(S_{k})_{k\geq 1}$ and $(T_{k})_{k\geq 0}$ defined recusrively by $T_{0}=0$ and for $k\geq 1$:
$R_{k}$ $:=\inf\{t\geq T_{k-1},\xi_{t}(o)=0\},$
$S_{k}$ $:=\inf\{t\geq R_{k},\xi_{t}(o)=1\},$
$T_{k}$ $:=\inf\{t\geq S_{k},H_{t}\leq 8\lambda\mathbbm{d}\}.$
We claim that:
  1. (i)
    Conditionally on $T_{k-1}$ being finite, the total time the center is infected on the time interval $[T_{k-1},T_{k}]$, which is bounded by $(R_{k}-T_{k-1})+(T_{k}-S_{k})$, is in average bounded by a constant, and more precisely
    $\operatorname{\mathbb{E}}\left[(R_{k}-T_{k-1})+(T_{k}-S_{k})\ \big{|}\ T_{k-1}<+\infty\right]\leq 1+16\log(1+\nicefrac{{9}}{{8}}).$
  2. (ii)
    Conditionally on $T_{k-1}$ being finite, we have a probability at least $\frac{1}{4}e^{-72\lambda^{2}\mathbbm{d}}$ of having $T_{k}$ infinite, and more precisely of failing to reinfect the center after time $R_{k}$, leading to $S_{k}$ being infinite.
Combining these two claims we deduce the result by
$\operatorname{\mathbb{E}}\left[\int_{0}^{\infty}\xi_{t}(o){\rm d}t\right]$ $\leq\sum_{k\geq 1}\mathbb{P}(T_{k-1}<+\infty)\operatorname{\mathbb{E}}\left[\int_{T_{k-1}}^{T_{k}}\xi_{t}(o)\rm dt\Big{|}T_{k-1}<+\infty\right]$
$\leq\left(1+16\log(1+\nicefrac{{9}}{{8}})\right)\cdot 4e^{72\lambda^{2}\mathbbm{d}}\leq 53e^{72\lambda^{2}\mathbbm{d}}.$
Thus it remains to check the claims in $(i)$ and $(ii)$.
For the first point, we clearly have $\operatorname{\mathbb{E}}[R_{k}-T_{k-1}|T_{k-1}<+\infty]\leq 1$, as the center recovers at rate 1 (if infected at time $T_{k-1}$). We continue with the following claim:
Claim.
Write $T_{[0,8\lambda\mathbbm{d}]}(H):=\inf\{t\geq 0,H_{t}\leq 8\lambda\mathbbm{d}\}$. Then, starting from any configuration with $H_{0}=h_{0}$, we have
$\operatorname{\mathbb{E}}\left[T_{[0,8\lambda\mathbbm{d}]}(H)\right]\leq 16\log\left(1+\frac{h_{0}}{8\lambda\mathbbm{d}}\right).$
Indeed, consider the process $Y_{t}=\log(1+H_{t}/8\lambda\mathbbm{d})+t/16$. For any $t<T=T_{[0,8\lambda\mathbbm{d}]}(H)$ and so $H_{t}>8\lambda\mathbbm{d}$, we can use the concavity of the log function and conditional Jensen inequality to see that the process $Y$ has nonpositive expected infinitesimal increase:
@@
$\frac{\operatorname{\mathbb{E}}\left[Y_{t+\rm dt}-Y_{t}|\mathcal{F}_{t}\right]}{{\rm d}t}\leq\frac{1}{8\lambda\mathbbm{d}}\cdot\frac{\lambda\mathbbm{d}-\frac{1}{4}H_{t}}{1+\frac{H_{t}}{8\lambda\mathbbm{d}}}+\frac{1}{16}=-\frac{1}{4}+\frac{1}{16}+\frac{3\lambda\mathbbm{d}}{8\lambda\mathbbm{d}+H_{t}}\leq 0,$
and therefore $Y_{t\wedge T}$ is a nonnegative supermartingale. We deduce
$\operatorname{\mathbb{E}}[T]\leq\operatorname{\mathbb{E}}[16Y_{T}]\leq\operatorname{\mathbb{E}}[16Y_{0}]=16\log\left(1+\frac{h_{0}}{8\lambda\mathbbm{d}}\right).$
It now follows that
$\begin{split}\operatorname{\mathbb{E}}[T_{k}-S_{k}|T_{k-1}<+\infty]&\leq 16\operatorname{\mathbb{E}}\left[\log\left(1+\frac{H_{S_{k}}}{8\lambda\mathbbm{d}}\right)\Big{|}T_{k-1}<+\infty\right]\\ &\leq 16\log\left(1+\frac{\operatorname{\mathbb{E}}[H_{S_{k}}|T_{k-1}<+\infty]}{8\lambda\mathbbm{d}}\right),\end{split}$
using again Jensen’s inequality. We thus have to control the conditional expectation $\operatorname{\mathbb{E}}[H_{S_{k}}|T_{k-1}<+\infty]$. To this end, observe that $T_{k-1}\leq 8\lambda\mathbbm{d}$ and that the expected infinitesimal increase of $H$ is bounded by $\lambda\mathbbm{d}$ by (9). The recovery of the root occurs at rate 1, and so
$\operatorname{\mathbb{E}}[H_{R_{k}}|T_{k-1}<+\infty]\leq 9\lambda\mathbbm{d}.$
Using again (9) when the center is healthy, we see that the process $(H_{t\wedge S_{k}})_{t\geq R_{k}}$ is a nonnegative supermartingale, which therefore has $\operatorname{\mathbb{E}}[H_{S_{k}}|T_{k-1}<+\infty]\leq 9\lambda\mathbbm{d}$. Note that we might have $S_{k}=+\infty$ if the recovery of the root is permanent, but in that case the result still holds (with $H_{S_{k}}=H_{\infty}=0$). Finally,
$\operatorname{\mathbb{E}}[T_{k}-S_{k}|T_{k-1}<+\infty]\leq 16\log(1+\nicefrac{{9}}{{8}}),$
with the understanding $T_{k}-S_{k}=0$ if $S_{k}=+\infty$. It remains to prove the second point.
Suppose that the centre is never reinfected after time $R_{k}$. If this happens, by (9) we see that $(e^{t/4}H_{t})_{t\geq R_{k}}$ is a nonnegative supermartingale. Hence by Markov’s inequality
$\mathbb{P}\left(\int_{R_{k}}^{\infty}H_{t}{\rm d}t>72\lambda\mathbbm{d}\bigg{|}T_{k-1}<+\infty\right)\leq\frac{9\lambda\mathbbm{d}\int_{R_{k}}^{\infty}e^{-(t-R_{k})/4}{\rm d}t}{72\lambda\mathbbm{d}}=\frac{1}{2}$
and so on the complement event, recalling that we infect the root at rate bounded by $\lambda H_{t}$, the probability that we really don’t ever see reinfection of the root is at least $\frac{1}{2}e^{-72\lambda^{2}\mathbbm{d}}.$ That is,
$\mathbb{P}\left(S_{k}=+\infty|T_{k-1}<+\infty\right)\geq\frac{1}{4}e^{-72\lambda^{2}\mathbbm{d}},$
which concludes the second point and thus the proof. ∎
@@
Consider $\left(\mathscr{T}_{\mathbbm{d},\left\lfloor\frac{1}{8\lambda^{2}}\right\rfloor},o\right)$ the infinite rooted tree with root $o$ of degree $\mathbbm{d}$ and all other vertices of degree $\left\lfloor\frac{1}{8\lambda^{2}}\right\rfloor$, with an infection sink $v$ at distance $\operatorname{d}(v,o)=k\geq 1$. The number of infections we expect it to receive from an infection at $o$ is
$m(o,v)\leq 27e^{72\lambda^{2}\mathbbm{d}}(2\lambda)^{k}.$